Perpetrators of a crime who have been given the maximum sentence of 10 years are believed by a fiscal to have been sentenced to no more than 7 years on average. A random sample of 20 of these crime offenders reveals a mean sentence of 7.6 years with a standard deviation of 2.4 years. Is the fiscal correct in his suspicions at 0.05 level of significance?


1
Expert's answer
2021-06-13T17:28:39-0400

"\mu=7 \\

n=20 \\

\bar{x}=7.6 \\

s=2.4 \\

α=0.05 \\

H_0: \mu=7 \\

H_1: \mu > 7 \\

df=n-1 = 20-1=19"

Test statistic:

"t= \frac{\bar{x}- \mu}{s / \sqrt{n}} \\

= \frac{7.6-7}{2.4 / \sqrt{20} } \\

= \frac{0.6}{0.5366} \\

= 1.118"

Read the t-table value at 5% level of significance at 19 degrees of freedom for Right tailed test is 1.72

t-cal<t-tab

So, we accept H0 and reject H1.

We can conclude, that the fiscal is correct in his suspicions at 0.05 level of significance.